# On the Geometry of Semantics in Next-token Prediction

### v1

@@ 3.1 NTP Objective as Sparse Soft-Label Classification
$\mathcal{L}({\bm{W}},\bm{\theta})=\sum\nolimits_{j\in[m]}\hat{\pi}_{j}\cdot\ell\left({\bm{W}}\bm{h}_{\bm{\theta}}(\bar{\bm{x}}_{j});\hat{\bm{p}}_{j}\right)\,,$ (1)
where the loss $\ell$ penalizes deviations between the model’s logits ${\bm{W}}\bm{h}_{\bm{\theta}}(\bar{\bm{x}}_{j})$ for context $j$ and its corresponding soft-label vector $\hat{\bm{p}}_{j}$. Unless otherwise specified, we use the standard cross-entropy loss we set the loss $\ell$ to be the standard CE loss, i.e., $\operatorname{CE}\left(\mathbb{S}({\bm{W}}\bm{h}_{\bm{\theta}}(\bar{\bm{x}}_{j}))\,||\,\bm{p}_{j}\right)$, where $\mathbb{S}()$ denotes the softmax map of logits to the probability simplex.
For later use, define the support matrix $\bm{S}\in\{0,1\}^{V\times m}$ of the conditional probability matrix ${\bm{P}}=[\hat{\bm{p}}_{1},\ldots,\hat{\bm{p}}_{m}]\in\mathbb{R}^{V\times m}$. Formally, $\bm{S}[z,j]=1\,\Leftrightarrow\,{\bm{P}}[z,j]>0$. For context $j$, we refer to tokens $z$ for which $\bm{S}[z,j]=1$ as in-support tokens, and refer to all others as off-support. Central to our analysis is the centered support matrix
$\widetilde{\bm{S}}:=(\mathds{I}_{V}-\frac{1}{V}\mathds{1}_{V}\mathds{1}_{V}^{\top})\bm{S}\,.$ (2)
For convenience, we refer to $\widetilde{\bm{S}}$ as the data-sparsity matrix, though note that unlike the support matrix $\bm{S}$, its entries are not binary due to the centering operation:
$\widetilde{\bm{S}}[z,j]=\begin{cases}1-\nicefrac{{|{\mathcal{S}}_{j}|}}{{V}}&\text{if $\bm{S}[z,j]=1$}\,,\\ -\nicefrac{{|{\mathcal{S}}_{j}|}}{{V}}&\text{if $\bm{S}[z,j]=0$}\,.\end{cases}\,$
3.2 Geometry of Words and Contexts
Following Yang et al. [2017], we assume the model is sufficiently expressive to allow optimizing context embeddings in (1) freely, instead of abiding by their architecture-specific parameterization. Concretely, this modeling abstraction yields a simplified training objective
$\min\nolimits_{{\bm{W}},{\bm{H}}}~{}~{}\mathcal{L}({\bm{W}}{\bm{H}})+\frac{{\lambda}}{2}\|{\bm{W}}\|^{2}+\frac{{\lambda}}{2}\|{\bm{H}}\|^{2}\,,$ (NTP-UFM)
which jointly optimizes ${\bm{W}}\in\mathbb{R}^{V\times d}$ and ${\bm{H}}:=[\bm{h}_{1},\ldots,\bm{h}_{m}]\in\mathbb{R}^{d\times m}$, where $\bm{h}_{j}$ is the freely optimized embedding of context $\bar{\bm{x}}_{j}$. Since the minimization is unconstrained for both variables, we follow Mixon et al. [2022], Wojtowytsch [2021], Fang et al. [2021] in referring to this as the unconstrained features model (UFM) for NTP training. We refer to ${\bm{W}}$ and ${\bm{H}}$ as the word and context embedding matrices.
Eq. (NTP-UFM) includes ridge-regularization with weight $\lambda>0$. Zhao et al. [2024] have analyzed the geometry of solutions to (NTP-UFM) when $\lambda\rightarrow 0$. This limit (referred to in the literature as the regularization-path) serves as a proxy for the limiting behavior of gradient descent (GD) training as the number of iterations approaches infinity Ji et al. [2020]. For large embedding dimensions $d\geq V$,  [^2] See Sec. A.1 for a discussion on this assumption. as $\lambda\rightarrow 0$, Zhao et al. [2024] show the following:
1. Logits Convergence: The logit matrix $\bm{L}={\bm{W}}{\bm{H}}$ decomposes into two orthogonal components: one that grows unboundedly during training and another that remains finite. After normalization (required to characterize the convergence of diverging logits), only the unbounded component becomes dominant, denoted as $\bm{L}^{\rm{mm}}$. Crucially, $\bm{L}^{\rm{mm}}$ depends only on the data support matrix $\bm{S}$ (see Zhao et al. [2024] for details).
2. SVD factors of $\bm{L}^{\rm{mm}}$: Like logits, word and context embeddings grow unboundedly in magnitude, but converge in direction to ${\bm{W}}^{\rm{mm}}=\bm{U}\bm{\Sigma}^{1/2}\mathbf{R}$ and ${\bm{H}}^{\rm{mm}}=\mathbf{R}^{\top}\bm{\Sigma}^{1/2}{\bm{V}}^{\top}$, respectively. Here, $\bm{U}\bm{\Sigma}{\bm{V}}^{\top}$ is the SVD of $\bm{L}^{\rm{mm}}$ and $\mathbf{R}$ is a partial orthogonal matrix.
3. Data-sparsity matrix as proxy: The matrix $\widetilde{\bm{S}}$ (see Eq. (2)) is a good proxy for $\bm{L}^{\rm{mm}}$.  [^3] This is formally proven in symmetric cases and supported by extensive empirical evidence. Put together, the singular factors of $\widetilde{\bm{S}}$ play a crucial role in determining the geometry of word and context embeddings learnt by NTP training. Denote this SVD as
$\widetilde{\bm{S}}:=\bm{U}\bm{\Sigma}{\bm{V}}^{\top},\quad$ (3)
where $\bm{U}\in\mathbb{R}^{V\times r}$, ${\bm{V}}\in\mathbb{R}^{m\times r}$ with $\bm{U}^{\top}\bm{U}={\bm{V}}^{\top}{\bm{V}}=\mathds{I}_{r}\,,$ the singular values $\bm{\Sigma}=\operatorname{diag}(\sigma_{1},\ldots,\sigma_{r})$ are ordered: $\sigma_{1}\geq\sigma_{2}\geq\ldots\geq\sigma_{r}>0\,$, and $r\leq V-1$ denotes the rank.
4 Geometry of Concepts
4.1 Motivating Questions
While prior work explicitly describes the geometry of word and context embeddings, it leaves open two key questions: (i) What is the latent information that drives the word/context embeddings to this geometry? And, (ii) does the resulting geometry form structures (e.g. in the form of clusters) that lend themselves to human-interpretable semantics?
To illustrate this, Fig. 2A shows a 3D projection of trained word and context embeddings on a simplified dataset for visualization. The conditional probability matrix ${\bm{P}}$ is shown on the right. The geometry reveals clear structure: animal-related words and contexts cluster on the left, plant-related ones on the right. This suggests the emergence of latent concept information—specifically, a positive/negative semantic distinction.
From Sec. 3.2, we know word embeddings take the form ${\bm{W}}=\bm{U}\sqrt{\bm{\Sigma}}\mathbf{R}$, where $\bm{U}$ encodes the row-space of $\widetilde{\bm{S}}$ (Eq. (2)). Since rows of $\widetilde{\bm{S}}$ correspond to words, and words that appear in similar contexts (e.g., “the organism that grows roots is…” vs. “the organism that has green leaves is…”) have similar patterns, $\bm{U}$ may naturally encode semantic relationships. Its columns can be interpreted as a “concept basis”, where a word’s projection reflects its association with these latent concepts.
4.2 Principal Components as Concepts
We term each dimension $k\in[r]$ resulting from the SVD of $\widetilde{\bm{S}}$ as a concept. Adopting terminology from Saxe et al. [2019], we call the columns $\mathbf{u}_{k}\in\mathbb{R}^{V}$ and $\mathbf{v}_{k}\in\mathbb{R}^{m}$ of matrices $\mathbf{U}$ and $\mathbf{V}$ the “word analyzer vectors” and “context analyzer vectors,” respectively. These vectors represent the alignment of tokens and contexts with each concept: for word $z$ and concept $k$, the sign of $\mathbf{u}_{k}[z]$ indicates association ($\mathbf{u}_{k}[z]>0$) or opposition ($\mathbf{u}_{k}[z]<0$) to the concept, while its magnitude quantifies the strength of this relationship.  [^4] Note that each word analyzer vector $\bm{u}_{k},k\in[r]$ has both positive and negative entries (corresponding to association and opposition), because $\bm{u}_{k}$ is orthogonal to $\mathds{1}_{V}$ (which is in the nullspace of $\widetilde{\bm{S}}$). Words for which $|\bm{u}_{k}[z]|$ is small are neutral to this concept. The same interpretation applies to context components $\mathbf{v}_{k}[j]$, measuring each context’s alignment with the respective concept. Fig. 1 verifies the SVD of $\widetilde{\bm{S}}$ captures the dataset’s semantic information for the simplified [^5] This design isolates semantic concept extraction by minimizing grammatical influences. These simplifications are for visualization clarity only—the principles apply to complex datasets with combined grammatical/semantic attributes and to full autoregressive training (see Secs. 5 and A.2). dataset with syntax "The organism that [attribute] is [subject]" (see Fig. 2).
To represent concepts in the embedding space and thus directly characterize their relationships with word and context embeddings, we define the $d$-dimensional “word-concept representations” $\bm{u}^{\rm{d}}_{k}$ and “context-concept representations” ${\bm{v}}^{\rm{d}}_{k}$ for $k\in[r]$ as projections onto the respective spaces of words and contexts, that is (see also Sec. B for details), $\bm{u}^{\rm{d}}_{k}={\bm{W}}^{\top}\bm{u}_{k}$ and ${\bm{v}}^{\rm{d}}_{k}={\bm{H}}{\bm{v}}_{k}.$ This ensures that words or contexts more closely aligned with a specific concept have embeddings closer to the concept’s representation. Similar to classical semantic axes formed by averaging relevant words, these representations are weighted averages where weights ($\bm{u}_{k}$) reflect concept relevance. Fig. 10 visualizes two most significant concept representations and their relationships with word/context embeddings.
Remark 1.
The idea that SVD factors of a co-occurrence-type matrix convey latent semantic information dates back at least to latent semantic analysis of word-document co-occurrence matrices Landauer [1997], Deerwester et al. [1990] and later to pointwise mutual information matrices of word-word co-occurrences Levy and Goldberg [2014], Levy et al. [2015]. Beyond SVD factorization, researchers have explored alternative factorization techniques, such as non-negative matrix factorization Lee and Seung [1999], Ding et al. [2006] and sparse dictionary learning Murphy et al. [2012], Faruqui et al. [2015], and have applied them to various transformations of co-occurrence data, including modifications that address the sparsity of PMI matrices by eliminating negative entries (e.g., PPMI) or applying probability smoothing techniques Levy et al. [2015]. Our key conceptual contribution is that we let the training dynamics NTP optimization determine both the matrix representation of the data and the factorization method used to encode latent semantics: the model naturally utilizes the centered support matrix $\widetilde{\bm{S}}$ and its SVD factorization. We validate through structured examples in this section and through experiments in Sec. 5, that the model’s emergent representational choices successfully encode linguistic components without explicit engineering of the representation space.
4.3 Orthant-based Clustering in Singular Spaces
Concepts extracted from $\widetilde{\bm{S}}$, represent significant components from data. However, as noted, for example, by Chersoni et al. [2021], Piantadosi et al. [2024], not all concepts individually correspond to linguistically interpretable factors. Instead, we posit that human-interpretable semantics can be identified by specific combinations of several concepts. Here, we formalize this hypothesis by relating it to an orthant-based clustering of word/context representations.
Recall from Sec. 4.2 that the $k$-th word analyzer vector $\bm{u}_{k}$ categorizes words $z$ based on $\texttt{sign}(\bm{u}_{k}[z])$: words with large $|u_{k}[z]|$ exhibit strong correlation with concept $k$, with $\texttt{sign}(\bm{u}_{k}[z])$ indicating whether this correlation is positive/negative. We extend this categorization beyond single dimensions to multiple concept dimensions simultaneously.
Let $\mathcal{K}=\{k_{1},\ldots,k_{p}\}$ denote a set of $p\leq r$ selected dimensions $k_{i}\in[r],i\in[p]$. Given $\mathcal{K}$, we define $2^{p}$ possible signature configurations $C=C(\mathcal{K})=[c_{k_{1}},\ldots,c_{k_{p}}]\in\{\pm 1\}^{p}$, where $c_{k_{i}}\in\{\pm 1\}$ indicates whether dimension $k_{i}$ contributes positively or negatively to the semantic category. Each signature configuration represents a potential semantic category, characterized by its member words and their degrees of association with the category.
Definition 1.
For $C=[c_{k_{1}},\ldots,c_{k_{p}}]$ with sign configurations $c_{k_{i}}\in\{\pm 1\}$ for dimensions $k_{i}\in[r]$:
  • A word $z$ is a member of $C$ (denoted $z\in C$) if and only if $\texttt{sign}(\bm{u}_{k_{i}}[z])=c_{k_{i}}$ for all $k_{i},i\in[p]$.
  • The typicality of a member $z\in C$ is defined as: $\text{Typicality}(z;C)=\sum\nolimits_{i\in[p]}\big{|}\bm{u}_{k_{i}}[z]\big{|}\,.$
Analogous definitions hold for contexts.
Geometrically, each configuration corresponds to a $p$-dimensional orthant (or orthant sheet) in the $d$-dimensional embedding space. Membership indicates which word/context embeddings lie within this orthant, effectively creating an orthant-based clustering in the subspace spanned by the selected concept dimensions $\mathcal{K}$. The typicality of a member measures the L1-norm (other choices are also possible) of its representation’s projection (the $z$-th row of $\bm{U}$ for words, $j$-th row of ${\bm{V}}$ for contexts) onto the selected concept dimensions. Fig. 2B and Sec. 5 demonstrate that this orthant-based clustering reveals interpretable semantic categories, while also inducing a hierarchical structure, where the number of concept dimensions determines the semantic granularity: broad linguistic categories (e.g., verbs) require fewer dimensions, while more specific categories (e.g., past-tense verbs) need more dimensions for precise characterization. See also Sec. C and Fig. 11 in the appendix.
4.4 Rate of Learning
We show that the hierarchical structure of semantic concepts discussed in Sec. 4.3 is reflected in their emergence during training: broad semantic distinctions are learned before finer-grained features. Specifically, we track the singular values of the logit matrix $\bm{L}_{t}={\bm{W}}_{t}{\bm{H}}_{t}$ at training iteration $t$ when minimizing the unconstrained (${\lambda}\rightarrow 0$) objective (NTP-UFM). On the one hand, at the end of training ($t\rightarrow\infty$), singular values exhibiting the largest divergence correspond to the largest singular values of $\widetilde{\bm{S}}$ Zhao et al. [2024], which represent the most important concepts. On the other hand, at initialization, all singular values of $\bm{L}_{t}$ are small (following typical small-scale initialization). As training progresses, Fig. 18 shows that the singular values increase at different rates: those that ultimately achieve the greatest magnitudes begin diverging earlier.
Fig. 3 provides empirical evidence of this hierarchical concept acquisition. The confusion matrices from a 2-layer transformer trained on natural language data reveal that the model first masters broad categorical distinctions (e.g., "Animals" vs. "Plants") before progressing to more specific semantic features. This progression hints at a coarse-to-fine learning order, which we formalize in the following theoretical analysis.
True ClassEpoch 0Refer to captionEpoch 20Refer to captionEpoch 40Refer to captionEpoch 200Refer to captionPredicted Class
Figure 3: Confusion matrices showing the progression of concept learning in a 2-layer transformer (embedding dimension $d=128$) across training epochs (0, 20, 40, 200). Each matrix cell shows the number of predictions for each true class. Classes correspond to support sets defined in Fig. 2C, with semantic labels (e.g., "Animals", "Plants") given in brackets where applicable. For example, the support set dog, wolf, cat, tiger, sparrow, eagle, trout, shark represents the "Animals" class (first row/column). The matrices illustrate a temporal learning hierarchy: broad categories like "Animals" and "Plants" emerge by epoch 20, while finer-grained concepts (e.g., "Reproduces with pollen", "Lives in water") develop later, between epochs 40–200. This reflects that high-variance (large singular value) concepts—often aligned with broader semantics—are learned earlier during training.
To gain theoretical insights into this behavior, we analyze (NTP-UFM) with square loss. [^6] This choice is motivated by the technical challenges in finite-time characterization of GD dynamics for CE loss due to coupling between singular directions. The simplified, square loss, could be seen as a first-order approximation of CE loss around zero initialization or in the high-temperature limit. Also, while CE loss is predominant in practice, we note that square-loss-type objectives have demonstrated practical success, as in the GloVe model Pennington et al. [2014], which achieves performance comparable to word2Vec Mikolov et al. [2013d]. We hope that future work can prove the analogous phenomena we observe empirically for CE loss. Specifically, we show in Sec. D in the appendix that the NTP-UFM setting is equivalent to the setting studied extensively by Saxe et al. [2013, 2019], Gidel et al. [2019]; thus, with sufficiently small initialization and learning rate, the $k$-th singular value of $\bm{L}_{t}$ undergoes a sharp phase transition from (almost) 00 to its limit $\sigma_{k}$ at time $T_{k}\propto 1/\sigma_{k}$ (see Thm. 1). In other words, more significant concepts (corresponding to larger singular values) are learned first, i.e. $T_{1}\leq T_{2}\leq\cdots\leq T_{r}$.
We further validate this behavior in a controlled imbalanced setting with one-hot labels. Two majority classes each appear across many contexts, while two minority classes appear in only one. As shown in Figure 9, the model initially predicts all inputs as the dominant majority class (Step 0), then begins separating the majority classes (Step 13), followed by gradual recognition of the minority classes (Step 30), and resolves confusion between them later in training (Step 55). This progression reflects the emergence of concepts in order of their singular values, consistent with theoretical analysis.
5 Experiments
5.1 Setup
Datasets. We conduct experiments on both synthetic data and pretrained models. For our synthetic experiments, we constructed two datasets by subsampling from two distinct corpora: TinyStories Eldan and Li [2023] and WikiText-2 Merity et al. [2017]. For both datasets, we fix the vocabulary size to $V=1,000$ and extract $m=10,000$ contexts, creating the "Simplified TinyStories/WikiText" datasets. We extracted the contexts by sampling frequent contexts of lengths 2-6 from the entire datasets. This variation in context-length emulates the autoregressive nature of training with sliding window, which yields rich data-sparsity patterns that in turn result in richer semantic information (compared to the last-token visuals in Sec. 4). For our experiments on pretrained transformers, we chose GPT-2 Radford et al. [2019] and extracted $50,000$ text sequences, each 11 tokens in length from GPT-2’s training corpus, the OpenWebText dataset. By segmenting these sequences into contexts of varying lengths (1-10 tokens), we generated $500,000$ contexts, which we designate as the "simplified OpenWebText" dataset.
Concepts and semantics. Following Sec. 4, for the simplified datasets where it is computationally feasible, we construct the $V\times m$ data-sparsity matrix $\widetilde{\bm{S}}$ and compute its SVD. The left/right singular vectors serve as our word/context concepts. As per Sec. 4.3, we also consider signature-based concept combinations, which we examine for semantic interpretability as discussed in Visuals below. For completeness, we also experiment with $k$-means spectral clustering as an alternative way to derive interpretable semantics from top important concepts. For the OpenWebText dataset, instead of performing SVD on $\widetilde{\bm{S}}$ (which is computationally expensive to construct and factorize), we work directly with the embeddings of the pretrained model. For word concepts, we extract ${\bm{W}}$ from GPT-2’s decoder and take the matrix $\bm{U}$ of word concepts to be its left singular matrix. For context concepts, we form a $d\times m$ matrix ${\bm{H}}$ by concatenating the embeddings (last-MLP layer representations before the decoder) of GPT-2 for each context in our subset. This matrix corresponds to only a portion of the “true” embedding matrix for the entire dataset. To ensure the coupling between ${\bm{W}}$ and ${\bm{H}}$ that is inherent to the SVD of $\widetilde{\bm{S}}$ (which we don’t have direct access to), we compute the matrix ${\bm{V}}$ of context concepts as ${\bm{V}}={\bm{H}}^{\top}\mathbf{R}\bm{\Sigma}^{-1/2}$, where $\bm{U}\bm{\Sigma}^{1/2}\mathbf{R}$ denotes the SVD of ${\bm{W}}$ that we already computed.

### v2

@@ 3.2 Interpretable Semantics via Orthant-based Clustering
This reveals a natural hierarchical structure in orthant-based clustering, where the number of concept dimensions determines semantic granularity: broad categories (e.g., verbs) require fewer dimensions, while specific categories (e.g., past-tense verbs) need more dimensions. See also Sec. 4.
3.3 Semantic Emergence
We now show that the semantics’ hierarchical structure is reflected in their emergence during training: broad categories are learned earlier than fine-grained ones. We formalize this by tracking the rate at which concepts are acquired throughout training. We find that concepts aligned with dominant singular directions (those corresponding to larger singular values) are learned first. Our analytic NTP-UFM model quantifies this claim as follows.
Theorem 1.
Consider gradient-descent minimization of NTP-UFM (Eq. (1)) with square loss. Assume infinitesimal step-size and spectral initialization ${\bm{W}}(0)=e^{-\delta}\bm{U}\mathbf{R}^{\top},{\bm{H}}(0)=e^{-\delta}\mathbf{R}{\bm{V}}^{\top}$ for partial orthogonal matrix $\mathbf{R}\in\mathbb{R}^{d\times r}$ ($\mathbf{R}^{\top}\mathbf{R}=\mathds{I}_{r}$) and initialization scale $e^{-\delta}$. Then, word and context embeddings at iteration $t$ evolve as ${\bm{W}}(t)=\bm{U}\sqrt{\bm{\Sigma}}\sqrt{{\bm{A}}(t)}\mathbf{R}^{\top}$ and ${\bm{H}}(t)=\mathbf{R}\sqrt{\bm{\Sigma}}\sqrt{{\bm{A}}(t)}{\bm{V}}^{\top}$ for ${\bm{A}}(t)=\operatorname{diag}(a_{1}(t),\ldots,a_{r}(t))$ with
Therefore, the $i$-th concept converges ($a_{i}(t)\rightarrow 1$) with exponential rate of $2\sigma_{i}$: concepts corresponding to larger singular values are learned first.
Proof and justification of the spectral initialization are in App. C.2. Our key insight is bridging the UFM with Saxe et al. [2013]’s study of closed-form learning-dynamics of two-layer linear nets.
To empirically verify this in natural language, we track concept learning through confusion matrices at different training timestamps. Since NTP performs soft-labeled classification where contexts can have multiple valid next-tokens, we define "effective classes" as the distinct sets of possible next-tokens (i.e., support sets) shared by at least one context in the dataset. For the dataset Fig. 2C, there are 9 such classes. Fig. 1C shows $9\times 9$ confusion matrices at four training timestamps from a 2-layer transformer; entry $(c,c^{\prime})$ indicates how many contexts from class $c$ are predicted to belong to class $c^{\prime}$. The diagonal entries show the number of correct classifications, with the total number of contexts per class shown after the slash. We observe that certain effective classes naturally correspond to semantic categories, as contexts sharing semantic meaning (e.g., describing animal attributes) tend to share the same support set (e.g., animal names). We annotate these associations with labels ("Animals," "Plants") in brackets on the confusion matrix axes.
The visualization reveals that the model first learns classes with coarse semantic associations (e.g., "Animals" vs. "Plants" at step 20), then progressively distinguishes classes with finer-grained semantic associations (e.g., "Lives in water (fish)" by step 40). This progression reflects that coarse semantic structures, corresponding to concepts with larger singular values, are learned earlier in training than finer-grained ones that require combining more concept dimensions.
To isolate the essential mechanism driving this progressive emergence, we identify the minimal data structure required: class imbalance alone suffices. Fig. 10 in Appendix demonstrates this on imbalanced MNIST, where the model learns distinctions in order of their corresponding singular values. Analyzer vectors with larger singular values encode majority-majority distinctions (learned first), the middle singular value encodes the majority-minority distinction (learned next), and smaller singular values encode minority-minority distinctions (learned last). This confirms that the progressive emergence of concept dimensions, ordered by singular value magnitude, stems from the fundamental optimization geometry determined by $\widetilde{\bm{S}}$ rather than being specific to linguistic structure. However, while even minimal imbalance produces meaningful categorical distinctions, richer sparsity patterns in $\bm{S}$, as naturally arise in the soft-labeled language setting, enable progressively richer semantic concepts to emerge from its SVD, as we show in Sec. 4.
4 Experimental Validation
Datasets. We construct datasets by subsampling from two distinct corpora: TinyStories [Eldan and Li, 2023] and WikiText-2 [Merity et al., 2017]. For both, we fix vocabulary size V=1V=1k and extract m=10m=10k contexts by sampling frequent sequences of lengths 2–6 from the original corpora, creating Simplified TinyStories and Simplified WikiText.
Models. To find semantics in models trained on controlled data, we train decoder-only transformers of varying sizes (6 or 12 layers, 1024 hidden dimensions) on Simplified datasets using AdamW until convergence (see Appendix D.2) for training details). We also analyze pretrained models starting with GPT-2-medium. To validate findings across architectures, scaling, and multi-language settings, we also analyze Qwen-7B (7B parameters with 4096-dimensional embeddings) [Bai et al., 2023]. Finally, while we focus primarily on NTP loss, the underlying structure of context-to-next-token co-occurrence is also present in masked language modeling (MLM) loss, where the context consists of surrounding tokens and the model is trained to predict randomly masked tokens within the sequence. To verify our analysis applies to here as well, we experiment with BERT-base [Devlin et al., 2018].
4.1 Methodology
Our approach differs depending on whether we have access to the training corpus. Simplified datasets: We explicitly construct $\widetilde{\bm{S}}$ and compute word and context analyzer vectors for each concept via SVD (Sec. 3.1). Pretrained models: We do not have access to their pretraining corpora, making it infeasible to construct $\widetilde{\bm{S}}$ and compute its SVD. We thus resort to an approximation: we extract the decoder matrix $\bm{W}$}, then use Arnoldi decomposition to compute their left singular vectors $\bm{U}$, which serve as proxies for word analyzer vectors.
For both cases, once we have analyzer vectors, we apply orthant-based clustering (Sec. 3.2). We combine the top-$p$ concepts $\{1,\ldots,p\}$ in successive order, starting from the most dominant. We find that choosing $p$ as small as 5 to 7 and inspecting word or context membership across different sign configurations (i.e., orthants) typically suffices to reveal interpretable semantic patterns.